What s graph of horizontal velocity and time when it is projected from ground

Grade:11

1 Answers

Shobhit
7 years ago
Q1:- 10 bullets per second mean 10gX10 is the rate of change of mass. Now F=vXdm/DT therefore F=100X10X10^-3 = 10NQ2:- impulse is equal to change in momentum or F.t=m( vf-vi)Now vf= 25. Vi=-30 therefore, f.t= 0.2(25-(-30))=0.2X85 =f.t therefore f=0.2X85/0.01 =1700N